### v1

@@ A.1.1 Bernoulli Bandit
A.1.2 Informative Action Bandit
In Section 6.2.2, we introduced an instantiation of information-directed sampling (IDS) (Russo & Van Roy, 2018) as an LLM agent design. The first bandit problem used for our evaluation is the contrived Example 2 of Russo & Van Roy (2018) meant to underscore one key deficiency in the exploration of TS. Here, we present the associated episodic regret curve of our experiment as further verification that our LLM-IDS behaves as one would expect of any IDS agent applied to the informative action bandit; namely, that exactly one sub-optimal but highly-informative selection of action 0 is needed before behaving optimally thereafter.
A.2 RiverSwim
@@
In Section 5.2.2, we reported positive results in a truncated (length-3) variant of the classic RiverSwim environment (Strehl & Littman, 2008) upon switching from GPT-4o to o1-mini as the underlying LLM for our PSRL implementation. For clarity, we use this section to detail the initial failures we encountered with GPT-4o in RiverSwim. Figure 15 shows the associated cumulative regret curves adhering to the same setup as outlined in Section 5.2.2, except we use $\kappa_{\pi^{\star}}=\kappa_{\mathrm{posterior}}=0$ and $\kappa_{\mathrm{sampling}}=0.5$. Despite achieving the best regret curve out of all presented LLM agents in RiverSwim, both of our LLM-based PSRL variants with GPT-4o incur near-linear regret while most instances of classic PSRL are able to achieve optimal behavior.
@@
We also report both vanilla and LLM-based PSRL run with prior distributions where all deterministic RiverSwim transitions (only those where the agent swims downstream) are given as prior knowledge. We posited that supplying all deterministic transitions as prior knowledge would fare better against classic PSRL. While this does allow LLM-based PSRL to exhibit optimal behavior in many trials, far too many still fail as the optimal policy LLM struggles to select optimal actions, even when supplied with posterior samples that have high fidelity to the true environment. Reasons for this include misread transition probabilities (such as swapping numerical values of the input posterior sample) as well as a lack of understanding for long-term planning. Additionally, we observe a rare occurrence where posterior updates can be prone to catastrophically forgetting a single transition, thereby halting learning progress entirely should the ommitted transition be essential to reaching the upstream reward.
Appendix B Experiment Prompts
@@
B.1 LLM-Based PSRL
@@
You are a Bayesian posterior distribution for a real-world sequential decision-making problem. Given a current prior belief about the environment and single trajectory observation, you should produce the posterior distribution that accurately reflects knowledge about possibly stochastic environment transitions and environment rewards based on the observed trajectory. A trajectory observation is a sequence of experiences, where each experience consists of a state, action, reward, and next state. Each unit of experience will be separated by XML $<$EXPERIENCE$>$ $<$/EXPERIENCE$>$ tags. The posterior distribution must always be complete and describe all sources of uncertainty the agent has about the world. There can be uncertainty about a stochastic transition or reward. The posterior distribution should take into account all information provided in the observed trajectory to update the prior belief about the environment. Be direct and don’t show your work. You cannot make any assumptions about the agent and the action selections used to generate the trajectory observation. Never try to model beliefs about the agent. Do not say anything beyond providing the posterior distribution. The agent’s interactions with the environment will generate rewards and the posterior distribution should keep track of how any and all rewards are generated. Information and knowledge in the current prior belief about the environment should never be discarded from the posterior distribution. If there is knowledge in the current prior belief about the environment that is unaffected by the trajectory observation, then this knowledge should not be changed and must be repeated exactly in the posterior distribution. Do not say anything to distinguish between old knowledge that is being retained and updated knowledge. The environment was described to the agent like this: <Environment Description>
Your current prior is as follows: <Input prior/LLM-generated posterior>. A trajectory observation is a sequence of experiences, where each experience consists of a state, action, reward, and next state. Each unit of experience will be separated by XML $<$EXPERIENCE$>$ $<$/EXPERIENCE$>$ tags. Here is an observed trajectory:<Full trajectory>. Remember that knowledge in the current prior must only be updated but can never be discarded, forgotten, or removed. Do not say anything about which information in the posterior is new and updated or old and remains the same from the prior.
@@
<Environment Description>. Always select optimal actions that maximize value across all future states and all remaining timesteps according to the following hypothesis: <LLM-generated posterior sample>. You must select actions that are optimal for and perfectly consistent with the above hypothesis. For each action, you must consider its immediate expect reward as well as the expected value of future states that can be visited by selecting the action. Always select from one of the available actions to take in the environment. Just say the action after ”Action: ” and nothing else.
@@
B.2 Multi-Armed Bandits
The environment description for the multi-armed bandit task was given as:
@@ B.2 Multi-Armed Bandits
You are a generator of Bernoulli bandit problems. A Bernoulli bandit problem is a collection of mean reward values, one for each available action. Knowledge about the reward of each available action will be given to you in the form of a Beta distribution representing beliefs about the mean reward of each arm. This knowledge will constrain the Bernoulli bandit problems you are allowed to generate. For each action, return one plausible hypothesis for the mean reward an agent will observe when taking that action. Each mean reward you return should be consistent with the knowledge you are given about the observed rewards of each action. Each action is independent and so each hypothesis you return for the mean reward of each action will be independent of all others. You must return real, numerical values starting with the phrase ”You think ” and do not say anything beyond providing the mean rewards of each action. You cannot just return the mean value of the Beta distribution as your guess for the mean reward. You must return a sample from each Beta distribution as your hypothesis. Before you return your mean reward values, describe how each one obeys all constraints and knowledge provided to you. The environment was described to the agent like this: <Environment Description>
@@
The environment description for the informative action multi-armed bandit was given as:

### v2

@@ B.1 Bernoulli Bandit
B.2 Customer Service Bandit & Prior (Mis)specification
To demonstrate one concrete instance of how our proposed LLM-based PSRL might meet the demands of a real-world decision-making problem, we adapt the customer service task of Tajwar et al. (2025) into a multi-armed bandit problem. In each of $K=20$ total time periods, the agent may either ask a question or offer a solution to address a customer issue randomly sampled from the dataset [^4] https://github.com/tajwarfahim/paprika/blob/main/llm_exploration/game/game_configs/customer_service.json of Tajwar et al. (2025). Similar to Tajwar et al. (2025), we use two additional LLMs to simulate the customer (who answers the agent’s questions and tries suggested solutions as a non-technical person would) and to be a judge/reward function who ultimately determines the binary reward indicating successful resolution of a customer’s issue. All models use GPT-4o as the underlying LLM.
For our LLM-based PSRL, we consider two methods for specifying the prior distribution that PSRL takes as input. In the first case, we simply ask GPT-4o to provide a prior distribution (a list of plausible underlying issues for the customer complaint as well as guessed probabilities based on how likely the model perceives the issue to be) that is given directly as input to our LLM-based PSRL agent. In preliminary experiments we found that, while this agent is capable of finding success often, it can suffer from issues of prior misspecification, where the true solution (also given in the dataset of Tajwar et al. (2025)) is not within the support of the LLM-generated input prior. To remedy this without giving away the solution, we use a second method of generating an input prior that guarantees it is well-specified; we provide the dataset solution for the sampled customer service issue to GPT-4o and indicate that it is one possible resolution but that GPT-4o must itself assign a probability to it based on how plausible it is perceived to be. We report the results of this latter agent as “well-specified” in Figure 5. All agents were run for a total of 20 trials.
Figure 12: Cumulative regret curves for the real-world customer service bandit task. All LLM agents use GPT-4o.
In the face of prior misspecification, something that the base PSRL algorithm does not entertain by assumption and therefore has no explicit mechanism to cope with, baseline LLM agent designs still cannot achieve a statistically significant improvement over PSRL. While theory is not a focus of this work, we simply note in passing that prior misspecification of posterior-sampling methods is a well-studied topic in bandit learning (Russo & Van Roy, 2014; Simchowitz et al., 2021; Liu et al., 2022), where one can provably expect a graceful degradation in performance commensurate with the degree of misspecification; colloquially, similar results are expected for the full RL setting as discussed, for instance, in the introduction of O’Donoghue (2021). Future work may greatly benefit from expanding on our results to more carefully examine how PSRL can remain robust in the face of such misspecified priors. Moreover, once the prior misspecification is removed (without handing the solution away as the agent must still sift through other plausible sources of customer issues), PSRL is able to demonstrate strong exploration that far exceeds baseline methods on a real-world task with a tremendously large action space.
Appendix C Early Failures with GPT-4o in RiverSwim
@@
In Section 4.2, we reported positive results in a truncated (length-3) variant of the classic RiverSwim environment (Strehl & Littman, 2008) upon switching from GPT-4o to o1-mini as the underlying LLM for our PSRL implementation. For clarity, we use this section to detail the initial failures we encountered with GPT-4o in RiverSwim. Figure 13 shows the associated cumulative regret curves adhering to the same setup as outlined in Section 4.2, except we use $\kappa_{\pi^{\star}}=\kappa_{\mathrm{posterior}}=0$ and $\kappa_{\mathrm{sampling}}=0.5$. Despite achieving the best regret curve out of all presented LLM agents in RiverSwim, both of our LLM-based PSRL variants with GPT-4o incur near-linear regret while most instances of classic PSRL are able to achieve optimal behavior.
@@
We also report both vanilla and LLM-based PSRL run with prior distributions where all deterministic RiverSwim transitions (only those where the agent swims downstream) are given as prior knowledge. We posited that supplying all deterministic transitions as prior knowledge would fare better against classic PSRL. While this does allow LLM-based PSRL to exhibit optimal behavior in many trials, far too many still fail as the optimal policy LLM struggles to select optimal actions, even when supplied with posterior samples that have high fidelity to the true environment. Reasons for this include misread transition probabilities (such as swapping numerical values of the input posterior sample) as well as a lack of understanding for long-term planning. Additionally, we observe a rare occurrence where posterior updates can be prone to catastrophically forgetting a single transition, thereby halting learning progress entirely should the omitted transition be essential to reaching the upstream reward.
Appendix D Limitation: Scaling Up Stochastic Environments
Appendix E Additional DeepSeek-R1 Results
While our experiments with RiverSwim (Figure 6) confirm the benefits of reasoning models that invest additional computational effort to produce so-called “reasoning” tokens prior to emitting response tokens, models such as o1-mini can be prohibitively expensive. To reduce these financial burdens and assess the efficacy of our proposed LLM-based PSRL with an alternative choice of constituent LLM, we present results for the combination lock (Figure 15 – 20 trials) and Wordle (Figure 8 – 40 trials) environments with DeepSeek-R1 (Guo et al., 2025).
Figure 15: Cumulative regret curves for the combination lock environment. Labels show the choice of constituent LLM model (GPT-4o or DeepSeek-R1) in each LLM agent.
Our results aggregated across both domains yield two key observations. At the highest level, we observe that R1 provides a performance improvement to all LLM agents (both ours and baselines). Curiously, we find that this performance improvement varies by model and domain; across both environments, we see very small improvements in Reflexion. Meanwhile, performance improvements for ICRL in the combination lock task and our LLM-based PSRL in Wordle are significant. More importantly, we find that the enhanced reasoning capabilities of DeepSeek-R1 applied to our best baseline LLM agents is not sufficient to yield a statistically-significant improvement over our proposed LLM-based PSRL, even when run with a “weaker” or less-capable GPT-4o as the constituent LLM. Such a result is somewhat reminiscent of classic boosting (Freund & Schapire, 1997), wherein an ensemble of weak learners are composed together into a strong (supervised) learner. Furthermore, these empirical results might (loosely) suggest that the strategic exploration strategy (specifically, Thompson Sampling) forged into the design and structure of the PSRL algorithm offers something beyond what a current strong reasoning model is capable of today, especially when given the freedom in action selections afforded by a LLM agent design like ICRL.
Appendix F Limitation: Beyond Thompson Sampling
While PSRL, through the use of TS, is known to yield a strong exploration strategy, it is by no means perfect. In the bandit literature, shortcomings of TS are well-known and naturally become more salient in the full RL problem (Russo & Van Roy, 2018; Lu et al., 2023). By only executing actions with some probability of being optimal, TS will never take sub-optimal actions that may yield tremendous information gain. Figure 3 already illustrates how a PSRL agent’s uncompromising execution of potentially-optimal policies cripples exploration and solely allows for the testing of two unknown letters at a time.
We offer two empirical evaluations to highlight the limitations of LLM-based PSRL exploration inherited from TS while also underscoring the future potential of our LLM-IDS. The first is a contrived but transparent multi-armed bandit problem given as Example 2 of Russo & Van Roy (2018). In this $(K+1)$-armed informative action bandit problem, there is a unique optimal action $A^{\star}\in[K]$ that yields a deterministic reward of 1 while all other arms yield a reward of 0; additionally, there is an action 0 that deterministically provides a reward equal to $(2\cdot A^{\star})^{-1}$. Naturally, an agent willing to deliberately select sub-optimal actions to gain information would take action 0 immediately and then produce optimal behavior thereafter with the identity of $A^{\star}$ in hand. Figures 17 and 17 show across 10 trials that LLM-IDS succeeds in recovering this optimal exploration strategy exactly for the $K=10$ instance whereas LLM-based PSRL is incapable of doing so while exploring via TS. This result also highlights one simple instance of the flexibility that specifying natural-language priors to LLM-based PSRL affords as encoding prior knowledge about the informative action might prove difficult when limited to classic statistical distributions. Extending past this contrived yet transparent bandit example, Figure 18 shows that LLM-IDS is able to outperform LLM-based PSRL in the combination lock task by more quickly testing for unknown digits while remaining unencumbered by known digits already discovered.
Appendix G Token Efficiency
In this section, we give a brief glimpse into the token efficiency of our proposed LLM-based PSRL agent relative to our two strongest baseline LLM agents, Reflexion and ICRL ($p=1.0$), using GPT-4o for all constituent LLMs. Notably, our focus in this work has been exclusively on data efficiency through prudent exploration and, as such, no concerted effort has been made in either our proposed agent or baseline agents towards optimizing for token efficiency explicitly (by selecting shorter prompts as inputs to the constituent LLMs) or implicitly (by encouraging LLMs to maintain brevity in their responses). With that said, Figures 20 and 20 illustrate token efficiency of these LLM agents in the combination lock and Wordle environments by plotting cumulative regret as a function of total tokens processed (on average).
Figure 19: Cumulative regret curves for the combination lock environment as a function of total tokens processed (on average).
Figure 20: Cumulative regret curves for the Wordle environment as a function of total tokens processed (on average).
In Figure 20, we see that, despite improved performance and actual convergence towards the optimal policy, our proposed PSRL-LLM consumes more tokens (on average) than Reflexion and ICRL in the combination lock environment. We suspect the primary driver behind the excess tokens comes from the tendency of GPT-4o to fully enumerate all possible correct codes in the “posterior” — something that neither baseline agent does thereby allowing them to be more economical with respect to the total number of tokens processed. Despite that, however, we see that our LLM-based PSRL does achieve better cumulative regret even if truncated to the same number of tokens processed by either baseline agent. In Figure 20, we see that LLM-based PSRL displays token efficiency that is comparable to Reflexion and superior to ICRL in Wordle, eventually able to more consistently identify target words in fewer turns than Reflexion, resulting in lower cumulative regret. Unlike in the combination lock environment, there are far too many possibilities for possible target words and GPT-4o never even attempts to enumerate these candidates, instead opting to maintain information about candidate correct letters and positions.
Appendix H Experiment Prompts
@@
Each LLM used in this work (both for our and baseline agents) was prompted to perform its designated function in the context of a broader agent design/algorithm (PSRL, Reflexion, ICRL, or ICPI). Thus, our prompt iteration process simply consisted of manually adjusting prompts until preliminary experiments showed the desired functionality being achieved. For baseline agents, especially those using ICL, this required few iterations; for some elements of PSRL that involve slightly more complicated entities than a policy; transition function; reward function; or evaluator, additional iterations were needed to weed out edge cases and tack on further constraints into the initial prompt used in the first iteration. For any given domain, the ability to successfully realize the desired functionality in each of the three LLMs should serve as “unit tests” signaling to an agent designer whether or not it is sensible to run our proposed PSRL agent. More generally, we make no claim that these prompts are optimal in any sense (a claim that likely no LLM agent paper can make in good faith). Investigating these choices in prompt iteration and downstream LLM agent robustness are important areas of future research.
H.1 LLM-Based PSRL
@@
You are a Bayesian posterior distribution for a real-world sequential decision-making problem. Given a current prior belief about the environment and single trajectory observation, you should produce the posterior distribution that accurately reflects knowledge about possibly stochastic environment transitions and environment rewards based on the observed trajectory. A trajectory observation is a sequence of experiences, where each experience consists of a state, action, reward, and next state. Each unit of experience will be separated by XML <<EXPERIENCE>> <</EXPERIENCE>> tags. The posterior distribution must always be complete and describe all sources of uncertainty the agent has about the world. There can be uncertainty about a stochastic transition or reward. The posterior distribution should take into account all information provided in the observed trajectory to update the prior belief about the environment. Be direct and don’t show your work. You cannot make any assumptions about the agent and the action selections used to generate the trajectory observation. Never try to model beliefs about the agent. Do not say anything beyond providing the posterior distribution. The agent’s interactions with the environment will generate rewards and the posterior distribution should keep track of how any and all rewards are generated. Information and knowledge in the current prior belief about the environment should never be discarded from the posterior distribution. If there is knowledge in the current prior belief about the environment that is unaffected by the trajectory observation, then this knowledge should not be changed and must be repeated exactly in the posterior distribution. Do not say anything to distinguish between old knowledge that is being retained and updated knowledge. The environment was described to the agent like this: <Environment Description>
Your current prior is as follows: <Input prior/LLM-generated posterior>. A trajectory observation is a sequence of experiences, where each experience consists of a state, action, reward, and next state. Each unit of experience will be separated by XML <<EXPERIENCE>> <</EXPERIENCE>> tags. Here is an observed trajectory:<Full trajectory>. Remember that knowledge in the current prior must only be updated but can never be discarded, forgotten, or removed. Do not say anything about which information in the posterior is new and updated or old and remains the same from the prior.
@@
<Environment Description>. Always select optimal actions that maximize value across all future states and all remaining timesteps according to the following hypothesis: <LLM-generated posterior sample>. You must select actions that are optimal for and perfectly consistent with the above hypothesis. For each action, you must consider its immediate expect reward as well as the expected value of future states that can be visited by selecting the action. Always select from one of the available actions to take in the environment. Just say the action after "Action: " and nothing else.
@@
H.2 Multi-Armed Bandits
H.2.1 Bernoulli Bandit
The environment description for the Bernoulli bandit task was given as:
@@ H.2.1 Bernoulli Bandit
You are a generator of Bernoulli bandit problems. A Bernoulli bandit problem is a collection of mean reward values, one for each available action. Knowledge about the reward of each available action will be given to you in the form of a Beta distribution representing beliefs about the mean reward of each arm. This knowledge will constrain the Bernoulli bandit problems you are allowed to generate. For each action, return one plausible hypothesis for the mean reward an agent will observe when taking that action. Each mean reward you return should be consistent with the knowledge you are given about the observed rewards of each action. Each action is independent and so each hypothesis you return for the mean reward of each action will be independent of all others. You must return real, numerical values starting with the phrase "You think " and do not say anything beyond providing the mean rewards of each action. You cannot just return the mean value of the Beta distribution as your guess for the mean reward. You must return a sample from each Beta distribution as your hypothesis. Before you return your mean reward values, describe how each one obeys all constraints and knowledge provided to you. The environment was described to the agent like this: <Environment Description>
@@
H.2.2 Customer Service Bandit
The environment description for the customer service bandit was given as:
You are going to role-play as a customer service agent and you have to help a customer resolve their issue. Your goal is to gather enough information to diagnose the problem and provide a correct solution. Your instructions are the following: 1.You may either ask the customer questions or suggest particular actions to the customer. 2. The customer may not be technically inclined, so keep your language simple and clear. 3.Avoid making assumptions — ask specific questions to determine the potential causes. You should guide the customer through basic troubleshooting steps and gather data on the situation. 4. You should try to make the customer satisfied and resolve their problem as quickly as possible. You should also keep your responses short and concise. 5. If the customer mentions a specific product they are using (for example, ABC electronics), then you are the customer support agent for that product/company, i.e., you represent that product or company and have to take appropriate actions without referring the customer to somewhere else. You will receive a reward of 1 if you succeed in resolving the customer’s issue and all other rewards are 0. The specific scenario the customer faces is this:<Troubleshooting task sampled from dataset>.
The posterior sampling LLM system prompts and user prompts were:
You are a troubleshooting hypothesis generator for a customer service agent. The initial issue faced by the customer was described to the customer service agent as follows: <Environment Description>. The customer service agent is trying to generate hypotheses for what the customer’s underlying issue really is. Given all knowledge the agent currently has currently obtained thus far about the customer’s issue, you must generate a single plausible hypothesis for what the customer issue is so the agent can correctly provide the solution to the customer. Current knowledge about the customer’s underlying issue will be given as a probability distribution listing possible underlying issues and the probability of those issues being accurate for the customer. The hypothesis you generate must be a sample from this distribution. While it is perfectly fine to return a sample that represents the element of the distribution with highest probability, you cannot just return the most likely hypothesis from this distribution simply because it has the highest probability. You must actually sample the distribution to generate your hypothesis. If the probabilities do not sum to 1 to form a valid probability distribution, sample a hypothesis based on what seems plausible using the knowledge available. Be as specific as possible when describing your hypothesis for the customer’s issue. You cannot just vaguely state that the customer’s item or some component of their item has an issue. You must be more precise than that. When you return your sample for the customer’s issue, start with the phrase "You think " and do not say anything beyond providing your hypothesis.
All current knowledge about the customer’s issue is as follows: <Input prior/LLM-generated posterior>. You must carefully read through this knowledge. Never say anything about the agent or tell the agent what to do.
Additionally, the respective system prompts used for the environment, judge/reward function, and prior-generator LLMs are as follows:
You are going to role-play as a customer experiencing a specific issue and interacting with a customer service agent. You must adhere to the following guidelines: 1. Only respond to the agent’s questions and provide relevant information when prompted. 2.Your responses should be concise and reflect a typical, non-technical customer’s knowledge level. 3. You are going to know what the solution for the specific issue is, but do not give away this information at any point. Your goal is to observe if the customer service agent can ask relevant questions and arrive at the specific issue. 4.When the agent asks you about a potential solution, you should determine what the outcome would be immediately (no asking the agent to wait or hold on) based on your knowledge about the true underlying problem, and let the agent know the result. For example, if your car is not starting because of lack of fuel, and the agent suggests you to check the fuel level in your car, it should fix your issue. Use your reasoning ability to decide whether the suggested solution works, given that the solution to your problem is: <Solution to sampled dataset issue>. 5. Reply with ’Goal reached’ (and nothing more) if the agent has reached or solved your particular issue. If your issue is solved, you must reply with ’Goal reached’. Your role-play starts now. The particular problem you, as the customer, are facing is: <Customer service issue sampled from dataset>, and the solution to this problem is: <Solution to sampled dataset issue>. Note that you should never give away or tell the agent what will solve your particular problem, your goal is to see if the agent can come up with the potential solution themselves by asking relevant questions and obtaining accurate information from you.
You are the judge of a customer service agent trying to help a customer in a particular scenario. You will be given the agent’s response, and you have to verify whether the agent has been able to successfully solve the problem of the customer correctly. You must adhere to the following guidelines: 1. If the agent provides a solution that is well aligned with the true solution, you must reply with <VALID>. 2. If the agent’s proposed solution does not align well with the true solution, then you must reply with <NOTVALID>. 3. The agent doesn’t need to match the solution word for word to be considered correct. The agent should be considered correct as long as their solution or question clearly demonstrates that the agent has correctly discovered the source of the customer’s issue. 4. Prior to returning your judgement of <VALID> or <NOTVALID> think about the agent solution and true solution and provide a brief justification of why they do or do not align well. The particular scenario the customer is facing is: <Customer service issue sampled from dataset>, and the true solution to their problem is: <Solution to sampled dataset issue>.
You are the generator of a prior distribution for a Bayesian decision-making agent. The agent is faced with a customer service task described as follows: <Customer service issue sampled from dataset>. The agent will be given an broad initial prior as follows: You know that rewards are binary and you will only receive a reward of 1 once the customer’s issue has been resolved. If you knew all the relevant details about the source of the customer’s issue, there would be no uncertainty about what correct solution to offer and obtain a reward of 1. You think that all common, reasonable issues based on the observations the customer has given are plausible. You think that more common and more realistic issues are more likely than uncommon and less realistic issues. Your job is to provide an additional supplement to this prior that is specific to the issue the customer is facing. Give a probability distribution for the possible underlying issue a customer could be faced along with the probabilities or relative likelihood for each issue you list based on which of them are more or less likely to be the culprit. (The next line is included if the prior is designed to be well-specified.) Be aware that one possible issue could be <Solution to sampled dataset issue> and include it in your prior with a probability the appropriately reflects how plausible it is to be the issue.
H.2.3 Informative Action Bandit
The environment description for the informative action bandit was given as: